# Aligning Cyber Space with Physical World: A Comprehensive Survey on Embodied AI

### v2

@@ VII-C Embodied Control
Embodied control aims to enable robots to acquire new skills through interaction and learning from their environment, thereby adapting to and completing complex tasks. Embodied control learns through interaction with the environment and optimizes behavior using a reward mechanism to obtain the optimal policy, thereby avoiding the drawbacks of traditional physical modeling methods. Embodied control methods can be divided into two types:
@@
Embodied control incorporates RL and sim2real techniques to optimize strategies through interaction with the environment, explore unknown domains, potentially surpass human capabilities, and adapt to unstructured environments. While robots can mimic many human behaviors, tasks often cannot be effectively completed without RL training based on environmental feedback. The most challenging scenes include contact-intensive tasks where manipulation requires real-time adjustments based on feedback such as the state, deformation, material, and force of manipulated objects, tasks where only RL proves adequate. In the era of MLMs, large models possess a generalized understanding of scene semantics, which can provide excellent reward functions for RL. Moreover, RL is also important for aligning large models. In the future, embodied agents, after pre-training and fine-tuning, still need RL to align with the physical world in order to better deploy themselves in real-world.
@@ VIII Challenges and Future Directions
Unified Embodied Foundation Model: Exploring foundation models for embodied robot tasks remains a nascent area of research, primarily due to the wide array of embodiments, environments, and tasks inherent in robotics. Compounding this challenge are isolated datasets and evaluation setups. Establishing a robust and unified foundation model for embodied robotics demands leveraging large-scale internet datasets and cutting-edge LLMs, MLMs and WMs.
Causal Relation Discovery: Existing data-driven embodied agents make decisions based on the intrinsic correlations within the data. However, this modeling approach does not allow the models to truly understand the causal relations between knowledge, behavior, and environment, resulting in biased strategies. This makes it difficult to ensure that they can operate in real-world environments in an interpretable, robust, and reliable manner. Therefore, it is important for embodied agents to construct embodied perception, reasoning, and interaction framework driven by world knowledge, capable of autonomous causal reasoning. By understanding the world through interaction and learning its workings via abductive reasoning [379], we can further enhance the adaptability, decision reliability, and generalization capabilities of multimodal embodied agents in complex real-world environments.
For embodied tasks (such as embodied question answering, visual language navigation, and instruction following), it is necessary to introduce embodied interactive causal representation learning. This involves establishing spatial-temporal causal relations across modalities through interactive instructions and state predictions, forming a representation learning system based on interaction and deduction. Moreover, agents need to understand the affordances of objects to achieve adaptive task planning and long-distance autonomous navigation in dynamic scenes. To optimize decision-making, it is necessary to combine counterfactual and causal intervention strategies [380, 381, 382, 383], trace causality from counterfactual and causal intervention perspectives, reduce exploration iterations, and optimize decisions. Constructing a causal graph based on world knowledge and driving sim-to-real transfer of agents through active causal reasoning will form a unified framework for embodied perception, reasoning, and interaction.

### v3

@@ VII-C Embodied Control
Embodied control learns through interaction with the environment and optimizes behavior using a reward mechanism to obtain the optimal policy, thereby avoiding the drawbacks of traditional physical modeling methods. Embodied control methods can be divided into two types:
@@
Embodied control integrates RL and sim-to-real techniques to optimize strategies through environmental interaction, enabling the exploration of unknown domains, potentially surpassing human capabilities, and adapting to unstructured environments. While robots can mimic many human behaviors, effective task completion often requires RL training based on environmental feedback. The most challenging scenarios include contact-intensive tasks, where manipulation necessitates real-time adjustments based on feedback such as the state, deformation, material, and force of manipulated objects. In such cases, RL is indispensable. In the era of MLMs, these models possess a generalized understanding of scene semantics, providing robust reward functions for RL. Additionally, RL is crucial for aligning large models with their intended tasks [400]. In the future, after pre-training and fine-tuning, RL is still required to align with the physical world, ensuring effective deployment in real-world environments.
@@ VIII Challenges and Future Directions
Causal Relation Discovery: Existing data-driven embodied agents make decisions based on the intrinsic correlations within the data. However, this modeling approach does not allow the models to truly understand the causal relations between knowledge, behavior, and environment, resulting in biased strategies. This makes it difficult to ensure that they can operate in real-world environments in an interpretable, robust, and reliable manner. Therefore, it is important for embodied agents to be driven by world knowledge, capable of autonomous causal reasoning. By understanding the world through interaction and learning its workings via abductive reasoning [402], we can further enhance the adaptability, decision reliability, and generalization capabilities of multimodal embodied agents in complex real-world environments.
For embodied tasks, it is necessary to introduce causal representation learning. This involves establishing spatial-temporal causal relations across modalities through interactive instructions and state predictions, forming a representation learning system based on interaction and deduction. Moreover, agents need to understand the affordances of objects to achieve adaptive task planning and long-distance autonomous navigation in dynamic scenes. To optimize decision-making, it is necessary to combine counterfactual and causal intervention strategies [403, 404, 405, 406], trace causality from counterfactual and causal intervention perspectives, reduce exploration iterations, and optimize decisions. Constructing a causal graph based on world knowledge and driving sim-to-real transfer of agents through active causal reasoning will form a unified framework for embodied perception, reasoning, and interaction.